OpenStudy (anonymous):

WILL MEDAL & FAN !!! HELP ASAP PLEASE !! Given triangle XYZ with angle bisector YW. The measurement of angle XYW is 2n + 10, the measure of angle WYZ is 4n – 34, and the measure of angles YZW and YXW are 58 and 62 degrees, respectively. Find the measure of angle XYZ. a. 108 b. 14 c. 32 d. 102

OpenStudy (anonymous):

@Nnesha

Nnesha (nnesha):

is there any diagram

OpenStudy (anonymous):

There appears to something wrong with your problem. My guess would be that you are suppose to find n for angle XYZ. Can you double check your problem?

OpenStudy (anonymous):

@marshallinwashington thats all that there was . there was no diagram .

OpenStudy (anonymous):

Well the sum of the angles in a triangle equals 180 degrees. You are given the value of two of these angles, YZW and YXW. So 58 + 62 = 120 degrees. So angle XYZ = 180-120= 60 degrees. However, 60 degrees is not one of choices listed in the answers. Now if the question was to find n for angle XYZ, you would have this equation to solve. 2n+10 +4n-34= 60 6n-24=60 6n=84 n=14

OpenStudy (anonymous):

Thankyou so much @marshallinwashington